# TxRay: Agentic Postmortem of Live Blockchain Attacks

### v1

@@ Abstract
To evaluate PoC correctness and quality, we develop PoCEvaluator, an independent agentic execution-and-review evaluator. On 114 incidents from DeFiHackLabs, TxRay produces an expert-aligned root-cause report and an executable PoC for 105 incidents, achieving 92.11 % end-to-end reproduction. Under PoCEvaluator, 98.1 % of TxRay PoCs avoid hard-coding attacker addresses, a +24.8pp lift over DeFiHackLabs, decoupling reproduction from attacker-controlled on-chain state and yielding self-contained, root-cause-consistent reconstructions rather than attacker-contract replays. In a live deployment, TxRay delivers validated root causes in 40 minutes and PoCs in 59 minutes at median latency. TxRay’s oracle-validated PoCs enable attack imitation, improving coverage by 15.6 % and 65.5 % over STING and APE.
@@ 1 Introduction
Postmortem correctness and quality evaluation: To measure DeFi PoC correctness and quality, we develop PoCEvaluator, an agentic execution-and-review evaluator: across 120 DeFiHackLabs incidents in a 426-day window (1 Oct 2024–30 Nov 2025, inclusive; after the gpt-5.1 training cutoff), we identify 114 ACT incidents and measure reproduction success (92.11 %) and two error rates: misalignment among ACT-identified cases (3.67 %) and miss rate (4.55 %). On the aligned subset (n=105), TxRay achieves 100% correctness, successfully compiling and passing runtime tests on a pinned on-chain fork. Quality metrics indicate self-contained reconstructions: no attacker-side artifacts appear in 104 / 105 cases; no real attacker addresses appear in 103 / 105 cases; and no exploit-specific constants appear in 53 / 105 cases. Under gpt-5.1 pricing, the estimated median API cost is $3.16 per run for root cause analysis and $1.25 per run for PoC generation.
@@ 2 Background
DeFi attacks, IDS, and postmortems. DeFi attacks are sequences of on-chain actions that violate protocol safety or liveness properties, e.g., by exploiting logic bugs, misconfigured parameters, or adversarial interactions between protocols. Many attacks are permissionless and publicly reproducible; in Section 3 we formalize ACT as the subset that an unprivileged adversary can realize using only public information and standard interfaces. When an exploit or anomalous behavior occurs, practice typically unfolds in two stages. First, intrusion-detection systems (IDS) (e.g., protocol-specific detectors, on-chain monitoring platforms, and incident dashboards) signal that an incident is in progress. Research and industry efforts focus heavily on this detection layer, from LLM-based anomaly detection [gai2023blockchain] to commercial systems such as BlockSec Phalcon [blocksec_dashboard], Hypernative [hypernative], and similar real-time monitoring services. Second, analysts perform postmortem reconstruction: recovering the precise sequence of on-chain actions, identifying the underlying vulnerability, and producing an executable PoC. To the best of our knowledge, this stage remains manual, leaving teams and the ecosystem to act under incomplete understanding: incidents may be misclassified as benign or non-exploitable, decisions may be wrong, and losses may continue until later analyses recover the true exploit mechanism and affected surface (e.g., the Balancer precision-loss exploit [balancer_attack_weilinli]).
@@ 4.2 Root Cause Analysis
(3) Root cause challenger checks the report draft against collected artifacts. It rejects drafts that omit lifecycle steps or roles, contradict traces/state diffs, or contain undetermined claims, and it requires an end-to-end causal account that states the violated invariant and the code-level breakpoint. On rejection, it returns structured feedback and missing-evidence requests; the orchestrator routes to re-analysis, re-collection, or seed-to-lifecycle expansion (Appendix A).
@@ 4.4 Implementation Details
Runtime environment. Our experiments run on an Ubuntu 22.04.5 server with an Intel i9-12900KS (16 cores / 24 hardware threads), 62 GiB RAM, and a 699 GiB NVMe system volume. Each run uses a session directory containing a .env file with credentials: (i) OPENAI KEY; (ii) ETHERSCAN KEY; and (iii) QUICKNODE KEY for archive RPC endpoints. At session start, the orchestrator copies JSON schemas and helper scripts, resolves chain-specific RPC URLs from a chainid-to-template map, and persists artifacts (traces, source or bytecode, JSON reports, Foundry projects, and logs). TxRay supports 22 EVM chains (Appendix C), requiring (i) archive+debug_traceTransaction RPC access (QuickNode), (ii) Etherscan v2 explorer APIs (account txlists, verified sources when available), and (iii) Foundry forked execution for PoC validation; when sources are unavailable, we use Heimdall as a bytecode-level fallback.
@@ 4.5 Walkthrough
Appendix A provides a compact end-to-end walkthrough with minimal iteration, giving intuition for the artifact flow and the orchestrator’s turn-by-turn control loop. In this case, our incident monitor extracts the seed transaction hash from a Twitter/X alert at 09:06 UTC on 1 Jan 2026; TxRay produces a validated root-cause report in 38 min and a validated PoC in 47 min from the alert. We highlight the Valinity walkthrough (Appendix B), which stress-tests TxRay under unverified attacker-side components and protocol-specific economic preconditions. The seed transaction calls an unverified contract, and the exploit hinges on wiring and preconditions not explicit in the seed trace. TxRay reconstructs an evidence-backed ACT narrative in which a flash-loan-driven swap sequence manipulates prices, triggers acquireByLTVDisparity() to mint VY and inflate the PAXG cap, and then reuses VY as collateral to open loans that unwrap and extract ETH-equivalent from WETH9 reserves. The following key takeaways summarize the main challenges in this walkthrough:
@@
(i) Filling information gaps under partial observability. Starting from a seed transaction, the full exploit context must be recovered, such as upgradeable wiring (proxyimplementation), registrar lookups, and per-asset configurations, and verified at the incident block. TxRay performs this autonomously by reconstructing , the on-chain pre-state in our ACT definition (Section 3.2), using targeted storage reads and state diffs rather than relying on inferred contract graphs. It then produces an ACT-compatible explanation of the transition . Result: In the Valinity case, closing these gaps required only four storage-slot queries and a pre-state diff trace for the seed transaction, successfully recovering the registrar wiring and per-asset configuration needed for an ACT-complete explanation in our Valinity walkthrough (Appendix B, Turns 3–6).
@@
(ii) Reasoning with imperfect and low-level artifacts. When attacker-side code is unavailable, analysts often rely on decompilation/disassembly (e.g., proprietary tools such as DeDaub) and debuggers (e.g., BlockSec Phalcon) that lift raw traces and state diffs into higher-level views. TxRay instead uses open-source bytecode tooling (less accurate than proprietary decompilers) and low-level RPC/explorer APIs, so it must lift evidence into higher-level claims while staying robust to imperfect decompilation. It grounds feasibility in txlists and opcode-level evidence (traces, disassembly, diffs), treating decompiler output as a hint and revising hypotheses when trace evidence contradicts it. Result: TxRay proceeded without verified attacker code by combining address-history queries, opcode-level traces (including a 14-transaction window around the incident block), and raw router bytecode/disassembly to justify unprivileged feasibility (Appendix B, Turns 2 and 6).
@@
(iii) Quality-driven iteration to reach reproducible artifacts. Human postmortems often stop at a plausible narrative and a “good-enough” reproducer (e.g., a shallow replay with hard-coded parameters). Our benchmark standard is higher: the final case must be executable, self-contained, parameter-free, and satisfy semantic oracles encoding the claimed mechanism. We enforce this with validator-driven rejection of non-conforming drafts; in our run, the reproducer required three attempts (two rounds of rejection and revision) before producing a self-contained Foundry PoC that re-creates the on-chain conditions (e.g., via a USDCPAXG Uniswap V3 swap) and satisfies the semantic oracles (Appendix B, Turns 7–9).
@@ 5.2 PoC Correctness and Quality Evaluation
PoCEvaluator Design. We implement PoCEvaluator to evaluate TxRay and DeFiHackLabs under a shared checklist (Table 1). Specifically, we follow CodeVisionary [wang2025codevisionaryagentbasedframeworkevaluating], which evaluates generated code via independent execution-based evaluation followed by multi-round negotiation and reports improved correlation with human judgments (+0.217 Pearson, +0.163 Spearman, +0.141 Kendall Tau over the best baseline on a 363-sample benchmark). PoCEvaluator instantiates this protocol for DeFi PoCs (Figure 3) in two stages. In stage (i) Agentic Independent Evaluation, it takes a PoC project directory and dispatches sub-evaluator agents to run forge test, inspect code and logs, and record execution artifacts (e.g., compilation status and pass/fail signals); sub-evaluators may also query on-chain context via Etherscan APIs, QuickNode RPC, and cast. In stage (ii) Multi-round Negotiation, an aggregator agent aggregates reports, isolates disagreements, and prompts sub-evaluators to maintain or revise decisions based on evidence until convergence, producing a consolidated report with decision histories (Appendix F).
@@
Quality Evaluation Result. On the aligned subset (n=105), Q1–Q3 measure whether a PoC reflects the inferred root-cause model and reproduces the strategy without relying on attacker-side artifacts, attacker account state, or attacker-designed constants (e.g., hard-coded calldata or exploit-specific numeric parameters). Under these metrics, TxRay passes Q1 in 104/105 cases (vs 102/105 for DeFiHackLabs), Q2 in 103/105 (vs 77/105), and Q3 in 53/105 (vs 14/105; Figure 4). Q4 (has_success_predicate) checks whether the PoC encodes validation oracles as assertions; TxRay passes Q4 in 105/105 aligned cases, and DeFiHackLabs passes Q4 in 3/105. Q5 (has_explanatory_comments) and Q6 (uses_address_labels) measure readability; TxRay passes Q5 in 93/105 cases (vs 63/105) and Q6 in 105/105 (vs 94/105).
@@ 6 Live Postmortem Pipeline
Postmortem Pipeline Implementation. Incident Monitor subscribes to a set of Twitter/X accounts (Appendix D). We seed this list with (i) incident announcement accounts and analysts referenced in DeFiHackLabs and (ii) additional accounts found via keyword search and manual screening. When a subscribed account posts, the monitor extracts txhash(es), resolves chain id(s), de-duplicates incidents, and uses gpt-oss-120b to filter for incident-related posts. For posts that pass, it forwards only txhash(es) and chain id(s) to TxRay (no narrative context), so TxRay reasons from on-chain evidence. We use Incident Monitor as an incident source and do not evaluate its coverage or precision. We seek to answer the following research questions:
@@ 9 Discussion and Limitations
Memorization. Throughout this paper, we evaluate TxRay on post-cutoff incidents (after the gpt-5.1 training cutoff, 30 Sep 2024) across the DeFiHackLabs benchmark, live postmortem pipeline, and incident dataset. The sole exception is generalized imitation (Section 7), which uses 1 Aug 2021–31 Jul 2022 because neither APE nor STING have open-sourced their code, preventing a post-cutoff comparison. To mitigate memorization risk, we inspect all reasoning traces to verify that TxRay derives exploits from on-chain evidence via explicit reasoning; readers should interpret imitation coverage results with this caveat.
Generalized Imitation vs. Mempool-time Execution. Section 7 shows that TxRay produces oracle-validated executable PoCs for the generalized imitation dataset, but its end-to-end latency is minutes (Figure 5 (right), Figure 7), far above an Ethereum block interval (12 s) and the same-block reaction budget. Most latency comes from model-facing reasoning (Figure 7), with additional overhead from on-chain data collection and third-party tooling (Figure 8). Online front-running requires either sub-block end-to-end latency or an execution environment that extends the reaction window by delaying inclusion of the adversary transaction. TxRay prioritizes evidence-grounded root-cause reporting and validation; imitation instead requires an executable value-capture strategy, not a human-readable narrative.
Dual use. Automated exploit reproduction is a double-edged sword: it can accelerate both researchers and attackers.
@@ 11 Conclusion
TxRay is an agentic postmortem system that reconstructs ACT exploits from seed transaction(s) by collecting public on-chain evidence and producing an evidence-backed root-cause explanation plus a deterministic, self-contained Foundry PoC. It grounds claims in traces/receipts/state diffs and accepts cases only when the PoC satisfies incident-specific semantic oracles, enabling standardized benchmarking and continuous curation of executable incident datasets. We view TxRay as a step toward faster, more reliable DeFi incident response.
Appendix A Walkthrough (PRXVTai Case)

### v3

@@ Abstract
To evaluate PoC correctness and quality, we develop PoCEvaluator, an independent agentic execution-and-review evaluator. On 114 incidents from DeFiHackLabs, TxRay produces an expert-aligned root-cause report and an executable PoC for 105 incidents, achieving 92.11 % end-to-end reproduction. Under PoCEvaluator, 98.1 % of TxRay PoCs avoid hard-coding attacker addresses, a +22.9pp lift over DeFiHackLabs, decoupling reproduction from attacker-controlled on-chain state (e.g., contracts deployed by the attacker, prerequisite approvals, or required funds) and yielding self-contained, root-cause-consistent reconstructions. In a live deployment, TxRay delivers validated root causes in 40 minutes and PoCs in 59 minutes at median latency. TxRay’s oracle-validated PoCs enable attack imitation, improving coverage by 15.6 % and 65.5 % over STING and APE.
@@ 1 Introduction
Postmortem correctness and quality evaluation: To measure DeFi PoC correctness and quality, we develop PoCEvaluator, an agentic execution-and-review evaluator: across 120 DeFiHackLabs incidents in a 426-day window (1 Oct 2024–30 Nov 2025, inclusive; after the gpt-5.1 training cutoff), we identify 114 ACT incidents and measure reproduction success (92.11 %) and two error rates: misalignment among ACT-identified cases (3.67 %) and miss rate (4.55 %). On the aligned subset (n=105), TxRay achieves 100% correctness, successfully compiling and passing runtime tests on a pinned on-chain fork. Quality metrics indicate self-contained reconstructions: no attacker-side artifacts appear in 104 / 105 cases; no real attacker addresses appear in 103 / 105 cases; and no exploit-specific constants appear in 53 / 105 cases. At the time of writing, our estimated median API cost is $3.16 per run for root cause analysis and $1.25 per run for PoC generation.
@@ 2 Background
DeFi attacks, IDS, and postmortems. DeFi attacks are sequences of on-chain actions that violate protocol safety or liveness properties, e.g., by exploiting logic bugs, misconfigured parameters, or adversarial interactions between protocols. Many attacks are permissionless and publicly reproducible; in Section 3 we formalize ACT as the subset that an unprivileged adversary can realize using only public information and standard interfaces. When an exploit or anomalous behavior occurs, practice typically unfolds in two stages. First, intrusion-detection systems (IDS) (e.g., protocol-specific detectors, on-chain monitoring platforms, and incident dashboards) signal that an incident is in progress. Research and industry efforts focus heavily on this detection layer, from LLM-based anomaly detection [gai2023blockchain] to commercial systems such as BlockSec Phalcon, Hypernative, and similar real-time monitoring services. Second, analysts perform postmortem reconstruction: recovering the precise sequence of on-chain actions, identifying the underlying vulnerability, and producing an executable PoC. To the best of our knowledge, this stage remains manual, leaving teams and the ecosystem to act under incomplete understanding: incidents may be misclassified as benign or non-exploitable, decisions may be wrong, and losses may continue until later analyses recover the true exploit mechanism and affected surface (e.g., the Balancer precision-loss exploit [balancer_attack_weilinli]).
@@ 4.2 Root Cause Analysis
(3) Root cause challenger checks the report draft against collected artifacts. It rejects drafts that omit lifecycle steps or roles, contradict traces/state diffs, or contain undetermined claims, and it requires an end-to-end causal account that states the violated invariant and the code-level breakpoint. On rejection, it returns structured feedback and missing-evidence requests; the orchestrator routes to re-analysis, re-collection, or seed-to-lifecycle expansion (Appendix C).
@@ 4.4 Implementation Details
Runtime environment. Our experiments run on an Ubuntu 22.04.5 server with an Intel i9-12900KS (16 cores / 24 hardware threads), 62 GiB RAM, and a 699 GiB NVMe system volume. Each run uses a session directory containing a .env file with credentials: (i) OPENAI KEY; (ii) ETHERSCAN KEY; and (iii) QUICKNODE KEY for archive RPC endpoints. At session start, the orchestrator copies JSON schemas and helper scripts, resolves chain-specific RPC URLs from a chainid-to-template map, and persists artifacts (traces, source or bytecode, JSON reports, Foundry projects, and logs). TxRay supports 22 EVM chains (Appendix E), requiring (i) archive+debug_traceTransaction RPC access (QuickNode), (ii) Etherscan v2 explorer APIs (account txlists, verified sources when available), and (iii) Foundry forked execution for PoC validation; when sources are unavailable, we use Heimdall as a bytecode-level fallback.
@@ 4.5 Walkthrough
Appendix C provides a compact end-to-end walkthrough with minimal iteration, giving intuition for the artifact flow and the orchestrator’s turn-by-turn control loop. In this case, our incident monitor extracts the seed transaction hash from a Twitter/X alert at 09:06 UTC on 1 Jan 2026; TxRay produces a validated root-cause report in 38 min and a validated PoC in 47 min from the alert. We highlight the Valinity walkthrough (Appendix D), which stress-tests TxRay under unverified attacker-side components and protocol-specific economic preconditions. The seed transaction calls an unverified contract, and the exploit hinges on wiring and preconditions not explicit in the seed trace. TxRay reconstructs an evidence-backed ACT narrative in which a flash-loan-driven swap sequence manipulates prices, triggers acquireByLTVDisparity() to mint VY and inflate the PAXG cap, and then reuses VY as collateral to open loans that unwrap and extract ETH-equivalent from WETH9 reserves. The following key takeaways summarize the main challenges in this walkthrough:
@@
(i) Filling information gaps under partial observability. Starting from a seed transaction, the full exploit context must be recovered, such as upgradeable wiring (proxyimplementation), registrar lookups, and per-asset configurations, and verified at the incident block. TxRay performs this autonomously by reconstructing , the on-chain pre-state in our ACT definition (Section 3.2), using targeted storage reads and state diffs rather than relying on inferred contract graphs. It then produces an ACT-compatible explanation of the transition . Result: In the Valinity case, closing these gaps required only four storage-slot queries and a pre-state diff trace for the seed transaction, successfully recovering the registrar wiring and per-asset configuration needed for an ACT-complete explanation in our Valinity walkthrough (Appendix D, Turns 3–6).
@@
(ii) Reasoning with imperfect and low-level artifacts. When attacker-side code is unavailable, analysts often rely on decompilation/disassembly (e.g., proprietary tools such as DeDaub) and debuggers (e.g., BlockSec Phalcon) that lift raw traces and state diffs into higher-level views. TxRay instead uses open-source bytecode tooling (less accurate than proprietary decompilers) and low-level RPC/explorer APIs, so it must lift evidence into higher-level claims while staying robust to imperfect decompilation. It grounds feasibility in txlists and opcode-level evidence (traces, disassembly, diffs), treating decompiler output as a hint and revising hypotheses when trace evidence contradicts it. Result: TxRay proceeded without verified attacker code by combining address-history queries, opcode-level traces (including a 14-transaction window around the incident block), and raw router bytecode/disassembly to justify unprivileged feasibility (Appendix D, Turns 2 and 6).
@@
(iii) Quality-driven iteration to reach reproducible artifacts. Human postmortems often stop at a plausible narrative and a “good-enough” reproducer (e.g., a shallow replay with hard-coded parameters). Our benchmark standard is higher: the final case must be executable, self-contained, parameter-free, and satisfy semantic oracles encoding the claimed mechanism. We enforce this with validator-driven rejection of non-conforming drafts; in our run, the reproducer required three attempts (two rounds of rejection and revision) before producing a self-contained Foundry PoC that re-creates the on-chain conditions (e.g., via a USDCPAXG Uniswap V3 swap) and satisfies the semantic oracles (Appendix D, Turns 7–9).
@@ 5.2 PoC Correctness and Quality Evaluation
PoCEvaluator Design. We implement PoCEvaluator to evaluate TxRay and DeFiHackLabs under a shared checklist (Table 1). Specifically, we follow CodeVisionary [wang2025codevisionaryagentbasedframeworkevaluating], which evaluates generated code via independent execution-based evaluation followed by multi-round negotiation and reports improved correlation with human judgments (+0.217 Pearson, +0.163 Spearman, +0.141 Kendall Tau over the best baseline on a 363-sample benchmark). PoCEvaluator instantiates this protocol for DeFi PoCs (Figure 3) in two stages. In stage (i) Agentic Independent Evaluation, it takes a PoC project directory and dispatches sub-evaluator agents to run forge test, inspect code and logs, and record execution artifacts (e.g., compilation status and pass/fail signals); sub-evaluators may also query on-chain context via Etherscan APIs, QuickNode RPC, and cast. In stage (ii) Multi-round Negotiation, an aggregator agent aggregates reports, isolates disagreements, and prompts sub-evaluators to maintain or revise decisions based on evidence until convergence, producing a consolidated report with decision histories.
@@
Quality Evaluation Result. Q1–Q3 in Table 1 evaluate whether a PoC reflects the inferred root cause and reproduces the strategy without relying on attacker-side artifacts, attacker account state, or attacker-designed constants (e.g., hard-coded calldata or exploit-specific numeric parameters). TxRay passes Q1/Q2/Q3 in 104/105, 103/105, and 53/105 cases, compared to 103/105, 79/105, and 17/105 for DeFiHackLabs (Figure 4). Q4 (has_success_predicate) measures whether the PoC encodes validation oracles as assertions: TxRay passes Q4 in 105/105 cases, while DeFiHackLabs passes Q4 in 3/105. Q5 (has_explanatory_comments) and Q6 (uses_address_labels) measure readability: TxRay passes Q5 in 93/105 and Q6 in 105/105, compared to 63/105 and 94/105 for DeFiHackLabs. Appendix A provides a case study.
@@ 6 Live Postmortem Pipeline
Postmortem Pipeline Implementation. Incident Monitor subscribes to a set of Twitter/X accounts (Appendix F). We seed this list with (i) incident announcement accounts and analysts referenced in DeFiHackLabs and (ii) additional accounts found via keyword search and manual screening. When a subscribed account posts, the monitor extracts txhash(es), resolves chain id(s), de-duplicates incidents, and uses gpt-oss-120b to filter for incident-related posts. For posts that pass, it forwards only txhash(es) and chain id(s) to TxRay (no narrative context), so TxRay reasons from on-chain evidence. We use Incident Monitor as an incident source and do not evaluate its coverage or precision. We seek to answer the following research questions:
@@ 9 Discussion and Limitations
Subjectivity. PoC quality metrics involve subjective judgment. Our checklist in Table 1 reflects one practical notion of quality, and the results should be interpreted accordingly rather than as a universal judgment of PoC style.
Memorization. We evaluate TxRay on post-cutoff incidents (after the gpt-5.1 training cutoff, 30 Sep 2024) across the benchmark, live postmortem pipeline, and incident dataset. The sole exception is generalized imitation (Section 7), which uses 1 Aug 2021–31 Jul 2022 because neither APE nor STING have open-sourced their code, preventing a post-cutoff comparison. To mitigate memorization risk, we inspect all reasoning traces to verify that TxRay derives exploits from on-chain evidence via explicit reasoning; readers should interpret imitation coverage results with this caveat.
Generalized Imitation vs. Mempool-time Execution. Section 7 shows that TxRay produces oracle-validated executable PoCs for the generalized imitation dataset, but its end-to-end latency is minutes (Figure 5 (right), Figure 7), far above an Ethereum block interval (12 s) and the same-block reaction budget. Most latency comes from model-facing reasoning (Figure 7), with additional overhead from on-chain data collection and third-party tooling (Figure 8). Online front-running requires either sub-block end-to-end latency or an execution environment that extends the reaction window. TxRay prioritizes evidence-grounded root-cause reporting and validation; imitation instead requires an executable value-capture strategy, not a human-readable narrative.
Dual use. Automated exploit reproduction is a double-edged sword. Refer to Section 12 for a detailed discussion.
@@ 11 Conclusion
TxRay is an agentic postmortem system that reconstructs ACT exploits from seed transaction(s) by collecting on-chain data and producing an evidence-backed root-cause explanation plus a self-contained Foundry PoC. It grounds claims in traces/receipts/state diffs and accepts cases only when the PoC satisfies oracles, enabling continuous benchmarking.
12 Ethical Considerations
This research is conducted with attention to ethical responsibilities, detailed below.
Controlled Fork-Based Assessment Environment. All postmortem analyses and PoC validations run in isolated environments and do not interact with production chains beyond read-only data collection. We validate synthesized PoCs by running Foundry tests against mainnet forks at the incident block on the 22 TxRay-supported EVM chains (Appendix E). These forks execute locally and deterministically, and the resulting transactions are never broadcast to a public mempool or included on-chain. We do not control every agent action, but we instruct the agent to use public data, and do not instruct the agent to interact with production blockchains by issuing transactions. We monitor agent runs and retain logs to audit behavior against these constraints; our review found no violations. Our evidence collection uses publicly available transaction data, traces, receipts/logs, and verified sources for contracts with verified code, and it does not involve private keys, privileged access, or any non-public user data. This design ensures that our experiments do not create financial externalities, do not manipulate real markets, and do not affect end users or protocol operations.
Stakeholders. The primary stakeholders are (i) protocol developers and security teams who need timely root-cause understanding, (ii) incident responders and analysts who require reproducible evidence, (iii) users and liquidity providers who seek faster containment and remediation, and (iv) the broader ecosystem of monitoring providers and researchers. A secondary stakeholder group includes adversaries who reuse or adapt research outputs for malicious purposes.
Dataset. Our standardized dataset is a curated record of historical incidents drawn from public sources (e.g., DeFiHackLabs[defihack_dashboard] and public security incident dashboards) and from our incident-monitor pipeline. The underlying evidence consists of public blockchain data such as transaction hashes, traces, logs, contract bytecode, and verified sources for contracts with verified code.
Privacy. We do not attempt to link on-chain addresses to off-chain identities, and our artifacts exclude off-chain personal identifiers. We treat blockchain addresses as sensitive identifiers because external information links addresses to individuals or organizations. To minimize privacy risks, our reports focus on technical mechanisms and on-chain evidence, avoid unnecessary attribution, and avoid including any off-chain personal details.
Disclosure for Newly Observed Incidents. For newly observed incidents from our live incident monitor, we follow a lagged disclosure policy to reduce risk. In particular, we publish a detailed root-cause report only after the protocol vulnerability and the relevant attack transactions have been publicly disclosed by security vendors. We delay public release of executable PoCs until the affected protocol has publicly acknowledged the issue and a remediation has been deployed, such as a contract upgrade or configuration change that removes the vulnerable behavior. When remediation status is unclear, we prioritize partial disclosure that supports measurement and reproducibility (e.g., incident metadata and high-level mechanism) without releasing a turnkey exploit reproduction.
Misuse Risk of Attack Imitation. Section 7 discusses that oracle-validated executable PoCs support attack imitation, including defensive actions such as reproducing the exploit to validate patches, censoring known-malicious transactions, recovering funds, and white-hat preemptive value extraction. Publishing higher-fidelity reconstructions increases misuse risk by lowering the effort required to reproduce observed incidents or adapt them to similar deployments. In its current implementation, TxRay is slower than APE [qin_usenix23] and STING [zhang2023your] and does not support transaction-time front-running. Evidence-backed root-cause reports and executable reproductions accelerate adversarial understanding of newly disclosed vulnerabilities and incident mechanisms, enabling imitation of newly observed attacks and rapid adaptation to similar deployments. This motivates controlled access to executable artifacts for newly observed incidents.
Artifact Availability and Potential for Misuse. Executable exploit reconstructions are subject to misuse. Our intent is defensive and scientific: to enable evidence-backed postmortems, measurable reconstruction quality, and standardized benchmarks that help the community evaluate detection, analysis, and response techniques. To reduce misuse risk while preserving reproducibility, we adopt a controlled artifact distribution strategy for components suited for operational weaponization, including executable incident reproductions for newly observed incidents. Access is granted under institutional verification and a responsible-use agreement, and we prioritize requests from researchers and defenders who demonstrate a legitimate need. We also design TxRay-generated PoCs to avoid relying on real attacker-controlled addresses and incident-specific attacker artifacts, reducing coupling to attacker infrastructure and emphasizing mechanism-level validation over replaying attacker deployments.
Appendix A PoC Quality Comparison (AAVEBoost)
To interpret our PoCEvaluator quality metrics (Q1–Q6 in Table 1), we use the AAVEBoost incident PoC 2025-06/AAVEBoost_exp.sol444https://x.com/CertiKAlert/status/1933011428157563188.. as a concrete case study. We compare the DeFiHackLabs PoC (as of paper submission) with the PoC generated by TxRay. Both PoCs satisfy correctness (C1–C3), but differ on quality metrics (Q1–Q6; Table 1). For each quality metric, we restate its checklist definition and compare the corresponding evaluation results for DeFiHackLabs and TxRay. To improve readability, we omit boilerplate code and truncate addresses. In the code boxes below, blue highlights indicate quality-metric passes and red highlights indicate failures.
DeFiHackLabs (✓). The PoC does not call the on-chain attacker contract; instead, it re-implements the attacker helper logic in-test (e.g., new AttackerC()).
TxRay (✓). The PoC avoids attacker-deployed contracts and drives the exploit through protocol entry points using deterministic helper roles (e.g., router).
DeFiHackLabs (). The PoC hard-codes the attacker EOA (0x5D44...) and executes the exploit under that identity (e.g., startPrank).
TxRay (✓). The PoC uses deterministic fresh addresses (e.g., makeAddr("attacker")), decoupling reproduction from attacker account state.
DeFiHackLabs (). The PoC hard-codes incident-specific numeric choices (e.g., 163, 3 * 10**17, and 48.9e18).
TxRay (✓). The PoC derives key quantities from on-chain state (e.g., REWARD()) and uses balance-based pre-checks and funding logic instead of fixed attacker-chosen constants.
DeFiHackLabs (). The PoC report does not assert profit or protocol state changes.
TxRay (✓). The PoC encodes success predicates as assertions (e.g., assertGt(attackerAfter, attackerBefore), assertLt(boostAfter, boostBefore)).
DeFiHackLabs (). Beyond header metadata, the PoC does not explain the loop bound, scaling factor, or funding amounts.
TxRay (✓). The PoC includes explanatory comments documenting preconditions and oracle intent (e.g., deterministic actors and the reward top-up rationale).
DeFiHackLabs (). The PoC uses raw addresses without role labels.
TxRay (✓). The PoC assigns role labels via vm.label (e.g., "AttackerEOA" and "RouterHelper").
Appendix B Key Prompts and Output Schema
We provide the key prompt templates and schemas used by TxRay and PoCEvaluator to support reproducibility. The full implementation is available in the artifact repositories.
Appendix C Walkthrough (PRXVTai Case)